OpenStudy (anonymous):

how do i solve this it is an open response for act practice our teacher gave us i do not understand 16x3y/4x+(5x-y)2

OpenStudy (anonymous):

solve what there's no = sign?

OpenStudy (anonymous):

i need to solve the whole problem lol sorry i forgot x equals -2 and y=3

razor99 (razor99):

i am here

razor99 (razor99):

now lets see your question

OpenStudy (anonymous):

ok held lol

razor99 (razor99):

um check out the answer in this site

OpenStudy (anonymous):

idk how lol it says to solve the equation and then i have to solve the other equation and see if they match ur way cuter than algebra lol

razor99 (razor99):

this one would be better

OpenStudy (anonymous):

ok im so confused lol

OpenStudy (anonymous):

its not helping lolz

razor99 (razor99):

just hold on okay plzz

razor99 (razor99):

i know the answer lol dont know the steps

OpenStudy (anonymous):

k lolz ll whats tha answer lolz i got 233 and i think that thats wrong

OpenStudy (anonymous):

lol hopeless uugh

razor99 (razor99):

thats the answer

OpenStudy (anonymous):

im still lost lol or maybe just lookin for an excuse to talk to u lolz

OpenStudy (anonymous):

ok well how do you solve (5x-y)2=25x2-10xy+y2when x is -2 and y is3

OpenStudy (anonymous):

im suppose to compare answers lol help please??

razor99 (razor99):

oh srry i had to go for a few minutes

OpenStudy (anonymous):

k well i have 15 minutes left lol check ur messages lolz

OpenStudy (anonymous):

how do i solve the last problem i added in a comment on this lolz

razor99 (razor99):

aww srry plzz i did not ignore u

OpenStudy (anonymous):

ok i wont be mad lolz ur really sweet for helping me on this lol

razor99 (razor99):

thanx

razor99 (razor99):

um may i have the permission to read your profile

razor99 (razor99):

hello u there

OpenStudy (anonymous):

u have my permission and why not ur to sweet to be single lol ok i need help for realz tho do u kno anything about distribuitive property??

OpenStudy (anonymous):

ok were working on solving multi step equations dioo u kno anything about how to solve them? ?

razor99 (razor99):

never heard of distribuitive property

OpenStudy (anonymous):

1/2(4x-6)-3x=7 how do i solve this??

OpenStudy (anonymous):

and like i said ur too cute and sweet and smart to be single ...lolz

razor99 (razor99):

oh we have to slove for equation x right

OpenStudy (anonymous):

right so how do i do that

OpenStudy (anonymous):

i think you multiply 1/2 by 4x and then by 6 right ?

razor99 (razor99):

yea

OpenStudy (anonymous):

ok i got 2x-3-3x=7 so now i combine x's right?

razor99 (razor99):

yea u are doin great :)

razor99 (razor99):

hello u there

OpenStudy (anonymous):

thnx ;) u made me blush lolz i got -5x-3=7

OpenStudy (anonymous):

idk if i did that right ??

razor99 (razor99):

hehehe wait i will show u how to do it now

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i think i got it backwards isnt it 1x -3=7 and then u add one to both sides of the equation ?

OpenStudy (anonymous):

idk what im doing wrong i keep getting that step confused somehow i thik its the positive and negatives

OpenStudy (anonymous):

think lol'

razor99 (razor99):